kristen's financial advisor has given her a list of 8 potential investments and has asked her to select and rank her favorite four. In how many different ways can she do this?
Answer by ikleyn(52781)

.

For any 8 given items, there are 8! = 8*7*6*5*4*3*2*1 = 40320 ways to order them (permutations).


You can place any of 8 item in the first position.

You can place any of 7 remaining items in the second position.

You can place any of 6 remaining items in the third  position.


   . . .  and so on . . . 


Multiplying these options, we obtain the number of permutations.


8 potential investments can be ranked in 8! = 8*7*6*5*4*3*2*1 = 40320 ways.

Solved, answered and explained.


------------------

This problem is about permutations.

On Permutations,  see introductory lessons
    - Introduction to Permutations
    - PROOF of the formula on the number of Permutations
    - Simple and simplest problems on permutations
    - Special type permutations problems
    - Problems on Permutations with restrictions

    - OVERVIEW of lessons on Permutations and Combinations
in this site.

Also,  you have this free of charge online textbook in ALGEBRA-II in this site
    - ALGEBRA-II - YOUR ONLINE TEXTBOOK.

The referred lessons are the part of this online textbook under the topic  "Combinatorics: Combinations and permutations".


Save the link to this textbook together with its description

Free of charge online textbook in ALGEBRA-II
https://www.algebra.com/algebra/homework/complex/ALGEBRA-II-YOUR-ONLINE-TEXTBOOK.lesson

into your archive and use when it is needed.